# SimQFL: A Quantum Federated Learning Simulator with Real-Time Visualization

### v1

@@ Abstract
Index Terms:
Quantum federated learning, quantum simulator, federated learning, simulator.
@@ I Introduction
Simulators are important in both conventional and quantum computing research because they enable algorithm development, testing, and validation in controlled and repeatable environments. In the field of federated learning (FL), simulators allow academics to experiment with collaborative training methodologies without the requirement for a large-scale distributed infrastructure. Open-source FL platforms, such as OpenFL [reina2021openfl] and PeerFL[luqman2024peerfl], have emerged as important tools, supporting a wide range of training scenarios and facilitating rigorous evaluation of federated algorithms across a variety of network, client, and data settings. Similarly, simulators such as Fujitsu’s 40-qubit system [imamura2022mpiqulacs] and platforms described at QSim 2024 [qsim2024conference] are critical environments for developing quantum circuits and evaluating quantum algorithms on classical hardware. These simulators not only decrease the barrier to testing but also provide insights into algorithm behavior, resource efficiency, and performance bottlenecks.
@@
Quantum machine learning (QML) investigates the integration of quantum computing principles with machine learning approaches in order to provide computational benefits such as faster processing, richer data representations, and improved model performance [khurana2024quantum]. Using quantum features such as superposition and entanglement, QML has brought new paradigms in data encoding, neural topologies, and optimization. However, one major problem with QML is that i) it requires data centralization, which creates data privacy concerns [pokharel2025quantum], and ii) centralized quantum processing, which requires significant quantum resources at the central server [innan2024fedqnn].
@@
By addressing these drawbacks with decentralized, privacy-preserving learning processes called federated learning (FL), quantum federated learning (QFL) shows significant potential as an extension to QML. Without disclosing raw data or needing centralized quantum processing, QFL allows many quantum clients to train local quantum models on confidential datasets and collaboratively update a shared global model [ren2023towards, rahman2025sporadic]. This configuration offers built-in benefits in data security, scalability, and error tolerance while significantly reducing the requirement for centralized quantum hardware. By modeling diverse client behaviors, multiple data distributions (including non-IID settings), and varied quantum configurations (such as the number of qubits or circuit depth), a well-designed QFL simulator can facilitate experimentation with these decentralized quantum models [qiao2024transitioning]. Such a simulator helps researchers evaluate performance trade-offs between privacy, accuracy, and utilization of resources, in addition to assisting with algorithmic development. However, current simulators exhibit two practical challenges.
@@
i) Lacking integrated support for QFL: QFL platforms lack quantum features like qubit encoding, variational quantum layers, or support for hybrid quantum-classical models, despite their great adaptability in conventional settings [park2025entanglement]. In contrast, general-purpose quantum simulators emphasize circuit reliability and quantum gates but provide minimal support for multi-client synchronization, model training, or optimization [gurung2025performance]. As a result, researchers studying QFL are frequently obliged to improvise tools or experiment with non-interactive scripts. This brings up the first important question: a) How can researchers use only classical hardware to efficiently build and examine the training dynamics of QML models in a federated environment?
@@
ii) Providing a visually appealing interface and real-time insight: In addition, current simulators are inadequate in terms of ease of execution and visually appealing during training. Subsequently, conventional FL or quantum simulation tools hardly enable real-time visualization, which is crucial for managing computing resources, detecting convergence, and troubleshooting quantum layers. For example, methods that investigate non-IID data impacts in QFL, such as the one by Zhao et al. [zhao2023non], depend on post-training assessments without ongoing observation. Comparably, privacy-preserving QFL frameworks that emphasize cryptographic integration, as MQFL by Dutta et al. [dutta2024mqfl], lack visual interpretability during training. This raises a second crucial question: b) How can researchers use visually appealing, real-time insights into performance and learning behavior to examine and refine QFL models interactively?
@@ II Related Works
Federated Learning (FL) has attracted a lot of interest for allowing group model training free from centralized data collection [rahman2025electrical, rahman2025multimodal, rahman2024multimodal]. Many open-source systems have been created to support FL research. Covering the top open-source FL frameworks, OpenFL [reina2021openfl] explores their respective benefits and drawbacks. PeerFL [luqman2024peerfl] also presents a simulator made for peer-to-peer FL at scale that integrates FL tools with the NS3 network simulator to replicate heterogeneous device experiments. These simulators provide the means to test FL algorithms and system behaviors in many contexts.
Essential tools for imitating quantum systems and enabling the evolution of quantum algorithms are quantum simulators. Using a CPU-based state vector approach to attain great computational efficiency, Fujitsu’s 40-qubit quantum simulator [imamura2022mpiqulacs] marks a major development in this field. Furthermore, providing a forum for debating the most recent advancements in quantum simulation, the QSim 2024 conference [qsim2024conference] brings together professionals from several fields to investigate both analog and digital methods.
To handle challenging computational tasks [pokharel2024electrocardiogram], QML investigates the integration of quantum computing concepts with machine learning methods. Providing a thorough overview of QML, including basic ideas, techniques, and statistical learning theories relevant to the subject, Wang and Liu [wang2024comprehensive]. Their work shows how, using quantum advantages, QML may transform data processing and analysis. Several architectures and approaches have been suggested in the quickly developing subject of QFL to improve the applicability and efficiency of distributed quantum machine learning. By providing a customizable and real-time visualizing platform for QFL experiments, the proposed SimQFL simulator aligns with and expands existing developments.
@@
Using genomic datasets on quantum simulators, Pokhrel et al. [pokhrel2024quantum] presented a data-encoding-driven QFL framework, showing its feasibility. Their work offers a proof of concept showing good results and highlights the need for data encoding in QFL. SimQFL expands on this by letting users choose data encoding techniques, therefore enabling a wider spectrum of applications.
In QFL, Zhao et al. [zhao2023non] investigated the difficulties of non-independent and identically distributed (non-IID) data. Using both theoretical and numerical studies, they suggested a structure to handle the non-IID problem. SimQFL uses procedures to replicate non-IID data distributions, therefore allowing researchers to investigate their effects on QFL models with relative efficiency.
Using quantum computing to improve performance while preserving data privacy via totally homomorphic encryption, Dutta et al. [dutta2024mqfl] presented a multimodal QFL framework. SimQFL enhances this by giving a forum for prototyping and analysis of privacy-preserving QFL models with real-time feedback.
Chehimi et al. [chehimi2023foundations] investigated the possibilities and difficulties of QFL over classical and quantum networks holistically. They pointed out important elements and special difficulties in implementing QFL, thereby providing fresh ideas and study paths. SimQFL provides an interactive and realistic environment for developing and tuning quantum neural networks in distributed contexts, therefore addressing some of these difficulties.
Emphasizing cooperative training without sharing local data, Yu et al. [yu2022quantum] presented a QFL architecture for distributed quantum networks. They developed a quantum gradient descent approach for local model training and techniques for data information extraction into quantum states. SimQFL extends this by offering adjustable quantum hyperparameters, such as qubits and quantum layers, enabling flexible experimentation with related collaborative training methodologies.
Diadamo et al. [diadamo2021qunetsim], designed to ease research and instructional activities at the quantum network level, is a Python-based quantum network simulator. It is fit for the testing and development of network-layer quantum protocols since it offers real-time simulation features and supports several quantum networking architectures. QuNetSim mostly concentrates on quantum communication and lacks built-in federated learning capabilities, even if it offers interactive real-time simulations. By combining real-time visualization and numerous customization choices, especially for QFL experiments, our suggested SimQFL explicitly expands such simulation ideas into QFL environments.
@@
QuNetSim [diadamo2021qunetsim] $\times$ Moderate Partial Real-time quantum network simulation with extensive quantum protocols
Apheris AI [reina2021openfl] $\times$ $\times$ Moderate $\times$ Comprehensive classical FL framework comparison
PeerFL [luqman2024peerfl] $\times$ Partial High Partial Peer-to-peer federated learning simulation at scale
Fujitsu Simulator [imamura2022mpiqulacs] $\times$ Partial Moderate Partial High-performance quantum state simulation
Pokhrel et al. [pokhrel2024quantum] $\times$ Moderate $\times$ Data-encoding-driven QFL using quantum simulators
Dutta et al. [dutta2024mqfl] $\times$ Limited $\times$ Multimodal QFL with homomorphic encryption
@@ IV-C Part 3: Selection of Hyperparameters
FL Setup: Users can select between two learning frameworks: Quantum Federated Learning (QFL) and Classical Federated Learning (FL). Although QFL is the default and most developed option in the current version, classical FL is supported and will be improved with new classical model choices. The simulator presently supports four commonly utilized benchmark datasets: MNIST [deng2012mnist], FashionMNIST [xiao2017fashion], and CIFAR-100 [krizhevsky2009learning]. These datasets offer varying levels of complexity and input dimensions, allowing users to evaluate their models under a variety of scenarios. The federated learning environment is mostly customizable. Users may select:

### v2

@@ I Introduction
Simulators are important in both conventional and quantum computing research because they enable algorithm development, testing, and validation in controlled and repeatable environments. In the field of federated learning (FL), simulators allow academics to experiment with collaborative training methodologies without the requirement for a large-scale distributed infrastructure. Open-source FL platforms, such as OpenFL [1] and PeerFL[2], have emerged as important tools, supporting a wide range of training scenarios and facilitating rigorous evaluation of federated algorithms across a variety of network, client, and data settings. Similarly, simulators such as Fujitsu’s 40-qubit system [3] and platforms described at QSim 2024 [4] are critical environments for developing quantum circuits and evaluating quantum algorithms on classical hardware. These simulators not only decrease the barrier to testing but also provide insights into algorithm behavior, resource efficiency, and performance bottlenecks.
@@
Quantum machine learning (QML) investigates the integration of quantum computing principles with machine learning approaches in order to provide computational benefits such as faster processing, richer data representations, and improved model performance [5]. Using quantum features such as superposition and entanglement, QML has brought new paradigms in data encoding, neural topologies, and optimization. However, one major problem with QML is that i) it requires data centralization, which creates data privacy concerns [6], and ii) centralized quantum processing, which requires significant quantum resources at the central server [7].
@@
By addressing these drawbacks with decentralized, privacy-preserving learning processes called federated learning (FL), quantum federated learning (QFL) shows significant potential as an extension to QML. Without disclosing raw data or needing centralized quantum processing, QFL allows many quantum clients to train local quantum models on confidential datasets and collaboratively update a shared global model [8, 9]. This configuration offers built-in benefits in data security, scalability, and error tolerance while significantly reducing the requirement for centralized quantum hardware. By modeling diverse client behaviors, multiple data distributions (including non-IID settings), and varied quantum configurations (such as the number of qubits or circuit depth), a well-designed QFL simulator can facilitate experimentation with these decentralized quantum models [10]. Such a simulator helps researchers evaluate performance trade-offs between privacy, accuracy, and utilization of resources, in addition to assisting with algorithmic development. However, current simulators exhibit two practical challenges.
@@
i) Lacking integrated support for QFL: QFL platforms lack quantum features like qubit encoding, variational quantum layers, or support for hybrid quantum-classical models, despite their great adaptability in conventional settings [11]. In contrast, general-purpose quantum simulators emphasize circuit reliability and quantum gates but provide minimal support for multi-client synchronization, model training, or optimization [12]. As a result, researchers studying QFL are frequently obliged to improvise tools or experiment with non-interactive scripts. This brings up the first important question: a) How can researchers use only classical hardware to efficiently build and examine the training dynamics of QML models in a federated environment?
@@
ii) Providing a visually appealing interface and real-time insight: In addition, current simulators are inadequate in terms of ease of execution and visually appealing during training. Subsequently, conventional FL or quantum simulation tools hardly enable real-time visualization, which is crucial for managing computing resources, detecting convergence, and troubleshooting quantum layers. For example, methods that investigate non-IID data impacts in QFL, such as the one by Zhao et al. [13], depend on post-training assessments without ongoing observation. Comparably, privacy-preserving QFL frameworks that emphasize cryptographic integration, as MQFL by Dutta et al. [14], lack visual interpretability during training. This raises a second crucial question: b) How can researchers use visually appealing, real-time insights into performance and learning behavior to examine and refine QFL models interactively?
@@ II Related Works
Federated Learning (FL) has attracted a lot of interest for allowing group model training free from centralized data collection [15, 16, 17]. Many open-source systems have been created to support FL research. Covering the top open-source FL frameworks, OpenFL [1] explores their respective benefits and drawbacks. PeerFL [2] also presents a simulator made for peer-to-peer FL at scale that integrates FL tools with the NS3 network simulator to replicate heterogeneous device experiments. These simulators provide the means to test FL algorithms and system behaviors in many contexts.
Essential tools for imitating quantum systems and enabling the evolution of quantum algorithms are quantum simulators. Using a CPU-based state vector approach to attain great computational efficiency, Fujitsu’s 40-qubit quantum simulator [3] marks a major development in this field. Furthermore, providing a forum for debating the most recent advancements in quantum simulation, the QSim 2024 conference [4] brings together professionals from several fields to investigate both analog and digital methods.
To handle challenging computational tasks [18], QML investigates the integration of quantum computing concepts with machine learning methods. Providing a thorough overview of QML, including basic ideas, techniques, and statistical learning theories relevant to the subject, Wang and Liu [19]. Their work shows how, using quantum advantages, QML may transform data processing and analysis. Several architectures and approaches have been suggested in the quickly developing subject of QFL to improve the applicability and efficiency of distributed quantum machine learning. By providing a customizable and real-time visualizing platform for QFL experiments, the proposed SimQFL simulator aligns with and expands existing developments.
@@
Using genomic datasets on quantum simulators, Pokhrel et al. [20] presented a data-encoding-driven QFL framework, showing its feasibility. Their work offers a proof of concept showing good results and highlights the need for data encoding in QFL. SimQFL expands on this by letting users choose data encoding techniques, therefore enabling a wider spectrum of applications.
In QFL, Zhao et al. [13] investigated the difficulties of non-independent and identically distributed (non-IID) data. Using both theoretical and numerical studies, they suggested a structure to handle the non-IID problem. SimQFL uses procedures to replicate non-IID data distributions, therefore allowing researchers to investigate their effects on QFL models with relative efficiency.
Using quantum computing to improve performance while preserving data privacy via totally homomorphic encryption, Dutta et al. [14] presented a multimodal QFL framework. SimQFL enhances this by giving a forum for prototyping and analysis of privacy-preserving QFL models with real-time feedback.
Chehimi et al. [21] investigated the possibilities and difficulties of QFL over classical and quantum networks holistically. They pointed out important elements and special difficulties in implementing QFL, thereby providing fresh ideas and study paths. SimQFL provides an interactive and realistic environment for developing and tuning quantum neural networks in distributed contexts, therefore addressing some of these difficulties.
Emphasizing cooperative training without sharing local data, Yu et al. [22] presented a QFL architecture for distributed quantum networks. They developed a quantum gradient descent approach for local model training and techniques for data information extraction into quantum states. SimQFL extends this by offering adjustable quantum hyperparameters, such as qubits and quantum layers, enabling flexible experimentation with related collaborative training methodologies.
Diadamo et al. [23], designed to ease research and instructional activities at the quantum network level, is a Python-based quantum network simulator. It is fit for the testing and development of network-layer quantum protocols since it offers real-time simulation features and supports several quantum networking architectures. QuNetSim mostly concentrates on quantum communication and lacks built-in federated learning capabilities, even if it offers interactive real-time simulations. By combining real-time visualization and numerous customization choices, especially for QFL experiments, our suggested SimQFL explicitly expands such simulation ideas into QFL environments.
@@
QuNetSim [23] $\times$ Moderate Partial Real-time quantum network simulation with extensive quantum protocols
Apheris AI [1] $\times$ $\times$ Moderate $\times$ Comprehensive classical FL framework comparison
PeerFL [2] $\times$ Partial High Partial Peer-to-peer federated learning simulation at scale
Fujitsu Simulator [3] $\times$ Partial Moderate Partial High-performance quantum state simulation
Pokhrel et al. [20] $\times$ Moderate $\times$ Data-encoding-driven QFL using quantum simulators
Dutta et al. [14] $\times$ Limited $\times$ Multimodal QFL with homomorphic encryption
@@ IV-C Part 3: Selection of Hyperparameters
FL Setup: Users can select between two learning frameworks: Quantum Federated Learning (QFL) and Classical Federated Learning (FL). Although QFL is the default and most developed option in the current version, classical FL is supported and will be improved with new classical model choices. The simulator presently supports four commonly utilized benchmark datasets: MNIST [24], FashionMNIST [25], and CIFAR-100 [26]. These datasets offer varying levels of complexity and input dimensions, allowing users to evaluate their models under a variety of scenarios. The federated learning environment is mostly customizable. Users may select: